A car starts from rest and moves with a constant acceleration for 8 sec .the net force is parallel to the displacement. What is the ratio of the increase in kinetic energy of the car over first and next 4sec interval of its motion?


1
Expert's answer
2021-05-07T17:27:11-0400

"v_0=0,~v_4=4a,~v_8=8a,"

"E_{40}=\frac m2(v_4^2-v_0^2)=8ma^2,"

"E_{84}=\frac m2(v_8^2-v_4^2)=24ma^2,"

"\frac{E_{84}}{E_{40}}=3."